# Longhorn: State Space Models are Amortized Online Learners

### v4

@@ 3.2 SSM as Online Learning and The Longhorn Architecture
$s_{t}=(I-\Delta_{t}k_{t}k_{t}^{\top})s_{t-1}+\Delta_{t}k_{t}x_{t},~{}~{}~{}\text{where}~{}\Delta_{t}=\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}\in[0,\infty).$ (7)
Proof.
As the objective in Equation 6 is in a quadratic form with respect to $s$, there is a unique minimum:
$\nabla L_{t}(s)=0$ $\Longleftrightarrow(s-s_{t-1})+\beta_{t}(s^{\top}k_{t}-x_{t})k_{t}=0$
$\Longleftrightarrow(I+\beta_{t}k_{t}k_{t}^{\top})s=s_{t-1}+\beta_{t}k_{t}x_{t}$
$\underbrace{\Longleftrightarrow}_{(3)}s=\bigg{(}I-\frac{\beta_{t}}{1+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}\beta_{t}k_{t}x_{t}$
$\Longleftrightarrow\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{(I+\beta_{t}k_{t}^{\top}k_{t}-\beta_{t}k_{t}k_{t}^{\top})\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
$\underbrace{\Longleftrightarrow}_{(5)}\bigg{(}I-\frac{\beta_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}k_{t}k_{t}^{\top}\bigg{)}s_{t-1}+\frac{\beta_{t}k_{t}x_{t}}{I+\beta_{t}k_{t}^{\top}k_{t}}$
(3) is derived from the fact that $(I+\beta_{t}k_{t}k_{t}^{\top})^{-1}=(I-\frac{\beta_{t}k_{t}k_{t}^{\top}}{I+\beta k_{t}^{\top}k_{t}})$ by the Sherman–Morrison formula. (5) is derived by noticing that $k_{t}^{\top}k_{t}k_{t}x_{t}-k_{t}k_{t}^{\top}k_{t}x_{t}=0$. ∎
@@
As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting a parallel scan. As a result, in practice, we use the diagonal approximation $1_{m}-\Delta_{t}k_{t}^{\odot 2}$ in place of $I-\Delta_{t}k_{t}k_{t}^{\top}$, where $x^{\odot 2}=x\odot x$ and $1_{m}$ is the $m$-dimensional vector of all ones. The final SSM recurrence therefore becomes:
$s_{t}=\underbrace{(1-\Delta_{t}k_{t}^{\odot 2})}_{A_{t}(x_{t})}\odot s_{t-1}+\underbrace{\Delta_{t}k_{t}x_{t}}_{B_{t}(x_{t})}.$ (8)
Following the Mamba design, when $x_{t}\in\mathbb{R}^{d}$ is a vector, we use $d$ parallel SSMs. As a result, $x_{t},\beta_{t},\Delta_{t}\in\mathbb{R}^{d},k_{t}\in\mathbb{R}^{m},S_{t}\in\mathbb{R}^{d\times m}$, Longhorn’s update becomes:
$S_{t}=A_{t}\odot S_{t-1}+B_{t},~{}~{}\text{where}~{}~{}A_{t}=(1-\Delta_{t}\otimes k_{t}^{\odot 2}),~{}~{}B_{t}=(\Delta_{t}\odot x_{t})\otimes k_{t}.$ (9)
@@
The final architecture of Longhorn follows Mamba strictly (Figure 3), and we only replace the SSM block with the Longhorn’s recurrence. We also provide an efficient CUDA kernel for it. The full inference-time algorithm is provided in Algorithm 1. One can compare Equation 9 to Equation 3 and other SSM in Section 2. Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forget gate is naturally derived from the key vector, i.e., $(1-k^{\odot 2}\otimes\Delta_{t})$.
Advantages of Longhorn: 1. While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem. 2. Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba). The forgetting is naturally linked to the key vector $k_{t}$ through the derivation. This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$. However, Longhorn demonstrates better performance even with fewer parameters than Mamba (See Figure 1 (left), Table 3, Table 4). 3. The closed-form solution in Equation 7 does not need any specific initialization. In contrast, Mamba requires careful initialization of the $A$ and $\Delta_{t}$. Compared to DeltaNet (Schlag et al., 2021; Yang et al., 2024), we restrict the state-transition matrix to be diagonal, hence Longhorn can leverage both parallel scan (as in Mamba) and the chunk-wise parallelization (as in GLA). Therefore, it is as fast as existing SSMs. 4. Unlike DeltaNet, which, as noted in Yang et al. (2024), cannot extrapolate beyond the training context, Longhorn demonstrates the capability to successfully extrapolate to contexts up to 16x longer at inference time (See Figure 1 (right)).
LA $\left\lVert s-s_{t-1}\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=s_{t-1}+k_{t}x_{t}$
RetNet $\gamma\left\lVert s-s_{t-1}\right\rVert^{2}+(1-\gamma)\left\lVert s\right\rVert^{2}-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\gamma s_{t-1}+k_{t}x_{t}$
GLA $(s-s_{t-1})^{\top}\mathrm{diag}(\alpha_{t})(s-s_{t-1})+s^{\top}\mathrm{diag}(1-\alpha_{t})s-2\langle s^{\top}k_{t},x_{t}\rangle$ $s_{t}=\mathrm{diag}(\alpha_{t})s_{t-1}+k_{t}x_{t}$
Longhorn $\left\lVert s-s_{t-1}\right\rVert^{2}+\beta_{t}\left\lVert s^{\top}k_{t}-x_{t}\right\rVert^{2}$ $s_{t}=(1-\Delta_{t}k_{t}^{\odot 2})s_{t-1}+\Delta_{t}k_{t}x_{t}$
$\Delta_{t}=\beta_{t}/(1+\beta_{t}k_{t}^{\top}k_{t})$
@@ 4 Related Work
This section provides a brief summary of recent advances in linear attention (LA) and deep state space models (SSMs).
@@
Several methods have been developed to address the quadratic complexity of the Transformer by making attention linear with respect to context length. In particular, Linformer uses a low-rank approximation of the self-attention by projecting the keys and values into a constant size matrix (instead of scaling with the sequence length)(Wang et al., 2020). Realizing that the main bottleneck causing the quadratic cost in Transformers is the Softmax function, the Linear Transformer replaces the Softmax function with a decomposable similarity function analogous to kernel methods, thereby making the computation linear with respect to sequence length(Katharopoulos et al., 2020). Performer approximates the softmax attention using positive orthogonal random features (Choromanski et al., 2020). More recently, based on the Linear Transformer, the Retentive Network (RetNet) adds additional constant forgetting and rotation (Sun et al., 2023). Gated Linear Attention (GLA) further experiments with learnable forget gates (Yang et al., 2023). Notably, linear attention can be viewed as a fast weight network where a slow net (the model’s parameters) learns to program a changing fast network (e.g., a linear predictor) by adapting its parameters (e.g., the $s_{t}$) online using inputs (Schlag et al., 2021).
@@
Instead of trying to linearize transformers, state space models (SSMs) start with parallelizable linear recurrent networks directly. Initially, the state transition matrix $A$ is assumed to be constant so that the recurrence can be computed in parallel using a convolution (Li et al., 2022; Gu et al., 2021). Subsequent developments include the Diagonal State Space (DSS) model (Gupta et al., 2022), Gated State Space (GSS) models (Mehta et al., 2022), S5 model (Smith et al., 2022), Bidirectional Gated SSM (BiGS)(Wang et al., 2022), H3 model(Fu et al., 2022), and Mamba (Gu & Dao, 2023). In addition, there are also works directly trying to make recurrent networks efficient, which often results in a particular form of SSM as well. This includes Deep Linear Recurrent Units (LRUs)(Orvieto et al., 2023; De et al., 2024), Hierarchically Gated Linear RNN (HGRN)(Qin et al., 2024b; a), and RWKV (Peng et al., 2023; 2024).
@@
The concept of a network modifying its own weights in response to inputs is not novel and has historical roots in the Fast-weight Programmer (Schmidhuber, 1992; 1993; Schlag & Schmidhuber, 2017; Schlag et al., 2021). Specifically, these models propose updating a weight matrix $W\in\mathbb{R}^{d\times m}$ using the outer product of two vectors, expressed as $\Delta W=v(x_{t})\otimes k(x_{t})$. This mechanism closely aligns with the principles underlying the Linear Attention model. Our framework builds upon the Fast Weight concept by tailoring the weight update process to address a specific online learning objective, thereby extending its applicability and effectiveness in dynamic learning environments.
@@
There are two concurrent works that share similar ideas to ours. Yang et al. (2024) propose a chunk-wise parallel form for scaling the DeltaNet (Schlag et al., 2021) to large-scale language modeling tasks (Yang et al., 2024). The update rule in DeltaNet can be viewed as conducting the gradient step instead of solving for the closed form of the online regression objective, hence the state transition matrix is $A(x_{t})=(I-\beta_{t}k_{t}k_{t}^{\top})$, which is not guaranteed to have eigenvalues $\leq 1$ (the forgetting can even be negative). Therefore, in practice Yang et al. (2024) normalize the key vector $k_{t}$ by its $\ell_{2}$ norm (Schlag et al., 2021), which can be restrictive. In contrast, the closed-form update in Longhorn ensures that the state transition matrix is always stable. In addition, $A(x_{t})$ is a matrix in DeltaNet while we apply a diagonal approximation ($k_{t}^{\odot 2}$). Therefore, Longhorn can leverage either parallel scan (as in Mamba) and chunk-wise parallel training (as in GLA), making it as fast as any existing SSM. In practice, we provide a parallel scan CUDA kernel, so Longhorn serves as a drop-in replacement for Mamba. On the other hand, Sun et al. (2024) propose the Test-Time Training framework, where the update of the state is derived from applying one gradient step of an online regression objective (Sun et al., 2024). To make the model parallelizable, the authors assume that each gradient step at $x_{t}$ ($t\geq 0$) is conducted at state $s_{0}$, hence deriving a parallel form that utilizes matrix multiplication. In contrast, Longhorn solves for the closed-form solution on every token, which potentially leads to a more flexible model.
@@ 5 Experiments
We conduct experiments on both synthetic benchmarks and the language modeling task to validate the performance of Longhorn:
  1. 1.
    To test Longhorn’s performance on memorizing associations, we compare it against other SSMs on the synthetic multi-query associative recall benchmark (Arora et al., 2023). We find that Longhorn is the only model that achieves near perfect recall rate when the sequence length scales up to 512 with a hidden dimension of 64).
  2. 2.
    To see how Longhorn performs on language modeling tasks, we first conduct the scaling law experiments using the OpenWebText dataset (Gokaslan & Cohen, 2019), with model sizes of 120M and 350M and context length 1024 or 4096, and then compare the validation perplexity of Longhorn against that of the other SSM models. We find that Longhorn consistently outperforms baseline deep SSM models across context lengths and model sizes.
  3. 3.
    In the end, to see how Longhorn performs on downstream language evaluation tasks, we train a 1.3B language model on the SlimPajama dataset (Soboleva et al., 2023) with 100B tokens, and then compare its downstream evaluation performance across 5 standard benchmarks. We find that Longhorn achieves 1.8x speedup sampling efficiency compared to Mamba, and also outperforms GLA and Mamba on average in downstream evaluation tasks.
@@ 5.1 Multi-Query Associative Recall
We first consider the synthetic benchmark Multi-Query Associative Recal (MQAR)l (Arora et al., 2023). The task is the agent will see a sequence of tokens $\{k_{1},v_{1},k_{2},v_{2},\dots,k_{T},v_{T}\}$, where each consecutive two-tokens become a key-value pair. At test time, the agent is provided with multiple $k\sim\{k_{1},\dots k_{T}\}$, the goal is for it to “retrieve” the corresponding values, hence the name multi-query associative recall. Following the original benchmark, we consider the sequence length $T\in\{64,128,256,512\}$ and model dimension (size of the latent embedding of a token) $d\in\{64,128,256,512\}$. We compare against 1) Transformer model (Attention), 2) Based architecture, which combines an SSM with local-attention, where the SSM is derived from the Taylor approximation of the self-attention (Arora et al., 2024), 3) Hyena (Poli et al., 2023), which is a special SSM that adopts long convolution via fast fourier transform, 4) RWKV (Peng et al., 2023), which can be viewed as the division of two SSMs (i.e., $y=a/b$, where $a,b$ are outputs from two SSMs). The state-transition matrix is a scalar, 5) BaseConv (Arora et al., 2023), an SSM that combines linear projection with convolution, and 6) Mamba (Gu & Dao, 2023), the state-of-the-art SSM that has data-dependent $A$ and $B$ (See Equation 3). Each experiment individually searches for the best learning rate from $\{1e-4,4.6e-4,2.2e-3,1e-02\}$ (evenly separated in the log space). Results are summarized in Figure 4.
Figure 4: Comparison of Longhorn against state-of-the-art SSM variants on the MQAR benchmark. The y-axis is the recall rate.
@@ 5.2 Scaling Law on OpenWebText
In this section we consider language modeling tasks on models with 120M or 350M parameters with 1024 or 4096 context length. We choose the OpenWebText dataset as it is small and serves as an easily accessible benchmark for quick benchmarks. [^2] We adapt the code from the nanoGPT repo, a minimal reproduction of the GPT-2 model. See https://github.com/karpathy/nanoGPT/tree/master. Following the Mamba (Gu & Dao, 2023) experiment setting, the training details are provided in Table 2:
125M 12 768 12 / 64 4800 6e-4 0.5M tokens 2.5B
350M 24 1024 16 / 64 13500 3e-4 0.5M tokens 7B
Table 2: Training details on OpenWebText.
We consider the following baseline models: LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), Mamba (Gu & Dao, 2023), RWKV (Peng et al., 2023), and GLA (Yang et al., 2023). Then we experiment with 1024 or 4096 context length $T$, which model sizes around 120M or 350M. Results are summarized in Figure 5 and Table 3.
@@
Table 3: Language modeling scaling law against LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), and Mamba (Gu & Dao, 2023). All models are trained on the OpenWebText dataset (Gokaslan & Cohen, 2019). Models vary from around 120M and 350M parameters, with context length 1024 or 4096.

### v5

@@ 3.2 The Longhorn Architecture
$S_{t,i}=(I-\varepsilon_{t,i}k_{t}k_{t}^{\top})S_{t-1,i}+\varepsilon_{t,i}k_{t}x_{t,i},~{}~{}~{}\text{where}~{}\varepsilon_{t,i}=\frac{\beta_{t,i}}{1+\beta_{t,i}k_{t}^{\top}k_{t}}\in[0,\infty).$ (6)
@@
Here, $S_{t,i}$ refers to the $i$-th row of $S_{t}$, $\beta_{t,i}$ refers to the $i$-th element of $\beta_{t}$. As $k_{t}k_{t}^{\top}$ is a matrix, it is hard to compute its cumulative product for conducting a parallel scan. As a result, in practice, we use the diagonal approximation $1_{m}-\varepsilon_{t,i}k_{t}^{\odot 2}$ in place of $I-\varepsilon_{t,i}k_{t}k_{t}^{\top}$, where $a^{\odot 2}=a\odot a$ and $1_{m}$ is the $m$-dimensional all-one vector. Following Mamba (Gu & Dao, 2023) and Transformer (Vaswani et al., 2017), we make $k_{t}=W_{k}x_{t}\in\mathbb{R}^{m}$ and $\beta_{t}=\sigma(W_{\beta}x_{t})\in\mathbb{R}^{d}$ (both are functions of $x_{t}$), where the activation $\sigma$ (the Sigmoid function) is to ensure that $\beta_{t}$ is positive and bounded. In summary, the final Longhorn update of $S_{t}$ becomes:
$S_{t}=A_{t}\odot S_{t-1}+B_{t},~{}~{}\text{where}~{}~{}A_{t}=(1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2}),~{}~{}B_{t}=(\varepsilon_{t}\odot x_{t})\otimes k_{t}.$ (7)
@@
The final architecture of Longhorn follows Mamba strictly (Figure 3), except that we replace the SSM block with Longhorn’s recurrence. We also provide an efficient CUDA kernel for it. The full inference-time algorithm is provided in Algorithm 1. One can compare Equation 7 to Equation 8 and other SSMs in Appendix A. Longhorn does not introduce an extra “forgetting” gate (hence it has fewer parameters), because the forgetting gate is naturally derived from the key vector, i.e., $(1_{d\times m}-\varepsilon_{t}\otimes k_{t}^{\odot 2})$.
Advantages of Longhorn
  1. 1.
    While we can derive the learning objective for some of the existing SSMs, Longhorn is the first SSM designed for explicitly solving an online regression problem.
  2. 2.
    Longhorn does not require a specific forget gate (e.g., $\alpha_{t}$ in GLA or $A$ matrix in Mamba). The forgetting is naturally linked to the key vector $k_{t}$ through the derivation. This saves about $\mathcal{O}(d\times m)$ parameters per SSM module, where $m$ is the dimension of $k_{t}$, and $d$ is the dimension of $x_{t}$. However, Longhorn demonstrates better performance even with fewer parameters than Mamba (See Figure 1 (left), Table 1, Table 2).
  3. 3.
    The closed-form solution in Equation 6 does not need any specific initialization. In contrast, Mamba requires a special careful initialization of the $A$ and $\varepsilon_{t}$.
  4. 4.
    Unlike DeltaNet (Yang et al., 2024), which struggles to extrapolate beyond training contexts, Longhorn successfully extrapolates to contexts 16x longer than it was trained for (Figure 1 (right)).
@@ 4 Related Work
This section provides a summary of recent advances in linear attention and state space models.
@@
Several methods reduce the quadratic complexity of Transformers by making attention linear with respect to context length. Linformer projects keys and values into a constant-size matrix, bypassing the scaling with sequence length (Wang et al., 2020). Linear Transformer replaces the Softmax function with a decomposable similarity function, achieving linear complexity (Katharopoulos et al., 2020). Performer approximates softmax attention using orthogonal random features (Choromanski et al., 2020). RetNet adds constant forgetting and rotation (Sun et al., 2023), while Gated Linear Attention introduces learnable forget gates (Yang et al., 2023). Linear attention can also be seen as a fast weight network where a slow net adapts a fast network’s parameters online using inputs (Schlag et al., 2021).
@@
State space models (SSMs) focus on parallelizable linear recurrent networks. Initially, a constant state transition matrix $A$ allows recurrence to be computed via convolution (Li et al., 2022; Gu et al., 2021). Key models include Diagonal State Space (DSS) (Gupta et al., 2022), Gated State Space (GSS) (Mehta et al., 2022), S5 (Smith et al., 2022), Bidirectional Gated SSM (BiGS) (Wang et al., 2022), H3 (Fu et al., 2022), and Mamba (Gu & Dao, 2023). Efficient recurrent networks often resemble SSMs, such as Deep Linear Recurrent Units (LRUs) (Orvieto et al., 2023; De et al., 2024), Hierarchically Gated Linear RNNs (HGRN) (Qin et al., 2024b; a), and RWKV (Peng et al., 2023; 2024).
@@
The idea of networks modifying their own weights in response to inputs dates back to the Fast-weight Programmer (Schmidhuber, 1992; 1993; Schlag & Schmidhuber, 2017; Schlag et al., 2021). These models update a weight matrix $W\in\mathbb{R}^{d\times m}$ via the outer product of two vectors: $\Delta W=x_{t}\otimes k(x_{t})$, a mechanism similar to Linear Attention. Our framework extends this concept by adapting the weight update process to suit specific online learning objectives, enhancing its use in dynamic learning environments.
@@
Two concurrent works share similar ideas with ours. Yang et al. (2024) propose a chunk-wise parallel approach to scale DeltaNet (Schlag et al., 2021) for large-scale language modeling. DeltaNet’s update rule, viewed as a gradient step for an online regression objective, results in a state transition matrix $A(x_{t})=(I-\beta_{t}k_{t}k_{t}^{\top})$, which can have eigenvalues $>$1, leading to instability. To address this, Yang et al. (2024) normalize the key vector $k_{t}$ by its $\ell_{2}$ norm, which can be restrictive. In contrast, Longhorn ensures stability with a closed-form update, using a diagonal approximation ($k_{t}^{\odot 2}$), allowing for both parallel scan (as in Mamba) and chunk-wise parallel training (as in GLA), making it as fast as existing SSMs. Additionally, we provide a parallel scan CUDA kernel, enabling Longhorn to serve as a drop-in replacement for Mamba. Sun et al. (2024) introduce the Test-Time Training framework, where state updates are derived from a gradient step on an online regression objective. To maintain parallelism, they assume each gradient step at $x_{t}$ uses the initial state $s_{0}$, enabling matrix multiplication. In contrast, Longhorn computes the closed-form solution for every token, offering greater flexibility.
@@ 5 Experiments
We validate Longhorn’s performance through the following experiments:
1)  We compare Longhorn against other SSMs on the multi-query associative recall benchmark (Arora et al., 2023) and find that Longhorn is the only model to achieve near-perfect recall at sequence lengths up to 512 with a hidden dimension of 64.
2)  Using the OpenWebText dataset (Gokaslan & Cohen, 2019), we assess Longhorn’s performance on language modeling with model sizes of 120M and 350M, and context lengths of 1024 or 4096, showing it consistently outperforms other SSMs in validation perplexity.
3)  We train a 1.3B language model on the SlimPajama dataset (Soboleva et al., 2023) with 100B tokens and compare its performance across 8 benchmarks, where Longhorn achieves better final performance and $>$1.8x better sample efficiency than Mamba and GLA.
4)  We additional apply Longhorn to vision domain and compare it against the Vision Mamba (ViM) (Zhu et al., 2024) model (Appendix 5.5), where Longhorn achieves performance comparable (slightly superior) to that of the ViM model.
@@ 5.1 Multi-Query Associative Recall
We first consider the synthetic benchmark Multi-Query Associative Recall (MQAR) (Arora et al., 2023). The agent observes a sequence of tokens $\{k_{1},v_{1},k_{2},v_{2},\dots,k_{T},v_{T}\}$, where each consecutive two-tokens become a key-value pair. At test time, the agent is provided with multiple $k\sim\{k_{1},\dots k_{T}\}$, the goal is to “retrieve” the corresponding values. Following the original benchmark, we consider the sequence length $T\in\{64,128,256,512\}$ and model dimension (size of the latent embedding of a token) $d\in\{64,128,256,512\}$. We compare against 1) Transformer model (Attention), 2) Based architecture, which combines an SSM with local-attention, where the SSM is derived from the Taylor approximation of the self-attention (Arora et al., 2024), 3) Hyena (Poli et al., 2023), which is a special SSM that adopts long convolution via fast fourier transform, 4) RWKV (Peng et al., 2023), which can be viewed as the division of two SSMs (i.e., $y=a/b$, where $a,b$ are outputs from two SSMs). The state-transition matrix is a scalar, 5) BaseConv (Arora et al., 2023), an SSM that combines linear projection with convolution, and 6) Mamba (Gu & Dao, 2023), the state-of-the-art SSM that has data-dependent $A$ and $B$ (See equation 8). Each experiment individually searches for the best learning rate from $\{10^{-4},4.6\times 10^{-4},2.2\times 10^{-3},10^{-2}\}$. Results are summarized in Figure 4.
Figure 4: Comparison of Longhorn to state-of-the-art SSMs on the MQAR benchmark. y-axis is the recall rate.
@@ 5.2 Scaling Law on OpenWebText
In this section, we consider language modeling tasks on models with 120M or 350M parameters with 1024 or 4096 context length. We choose the OpenWebText dataset as it is small and serves as an easily accessible benchmark for quick benchmarks. [^1] We adapted code from the nanoGPT repository https://github.com/karpathy/nanoGPT, which is a minimal reproduction of GPT-2 model using PyTorch. The details about the architecture is provided in Appendix D. We consider the following baseline models: LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), Mamba (Gu & Dao, 2023), RWKV (Peng et al., 2023), and GLA (Yang et al., 2023). Then we experiment with 1024 or 4096 context length $T$ and model sizes around 120M or 350M. Results are summarized in Table 1 and Figure 5.
@@
Table 1: Language modeling scaling law against LLaMA (Touvron et al., 2023), RetNet (Sun et al., 2023), RWKV (Peng et al., 2023), and Mamba (Gu & Dao, 2023). All models are trained on the OpenWebText dataset (Gokaslan & Cohen, 2019). Models vary from 120-350M parameters and 1024-4096 context length.